# CatCMA with Margin: Stochastic Optimization for Continuous, Integer, and Categorical Variables

### v5

@@ Appendix A Procedure of Modified Margin Correction
(52) $\exists i\in\{1\!:\!\lambda,\ldots,\mu\!:\!\lambda\}\enspace\text{s.t.}\enspace[\textsc{Enc}(\boldsymbol{x}_{i})]_{j^{\mathrm{in}}_{n}}\neq[\textsc{Enc}(\boldsymbol{m}^{(t)})]_{j^{\mathrm{in}}_{n}}$
For each $n\in\{1,\ldots,{N_{\mathrm{in}}}\}$, the margin correction updates $[\boldsymbol{m}^{(t+1)}{\mskip-2.0mu}]_{j^{\mathrm{in}}_{n}}$ and $\langle\boldsymbol{A}^{(t)}\rangle_{j^{\mathrm{in}}_{n}}$ according to whether $n$ is the dimension in which the successful integer mutation occurred.
Case of ${}[\textnormal{{Enc}}(\boldsymbol{m}^{(t+1)})]_{j^{\mathrm{in}}_{n}}\in\{z_{n,1},z_{n,L_{n}}\}$.
@@
(54) $p^{({t+1})}_{\mathrm{mut},{n}}\leftarrow\max\left\{\alpha,p^{({t+1})}_{\mathrm{mut},{n}}\right\}\enspace.$
@@
(55) $p^{({t+1})}_{\mathrm{mut},{n}}\leftarrow\max\left\{\alpha,\min\left\{p^{({t+1})}_{\mathrm{mut},{n}},p^{({t})}_{\mathrm{mut},{n}}\right\}\right\}\enspace.$
@@
Finally, we correct $[\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}$ as follows:
@@
Case of ${}[\textnormal{{Enc}}(\boldsymbol{m}^{(t+1)})]_{j^{\mathrm{in}}_{n}}\in\{z_{n,2},\ldots,z_{n,L_{n}-1}\}$.
Let us denote the nearest two thresholds to $[\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}$ as follows:
@@
(58) $p^{({t+1})}_{\mathrm{low},{n}}$ $=\Pr\left([\boldsymbol{v}]_{j^{\mathrm{in}}_{n}}\leq\ell^{({t+1})}_{\mathrm{low},{n}}\right)\enspace,$
(59) $p^{({t+1})}_{\mathrm{up},{n}}$ $=\Pr\left(\ell^{({t+1})}_{\mathrm{up},{n}}<[\boldsymbol{v}]_{j^{\mathrm{in}}_{n}}\right)\enspace,$
(60) $p^{({t+1})}_{\mathrm{mid},{n}}$ $=1-p^{({t+1})}_{\mathrm{low},{n}}-p^{({t+1})}_{\mathrm{up},{n}}\enspace.$
@@
(72) $p^{({t+1})}_{\mathrm{mut},{n}}\leftarrow p^{({t+1})}_{\mathrm{low},{n}}+p^{({t+1})}_{\mathrm{up},{n}}\enspace.$
Finally, $[\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}$ and $\langle\boldsymbol{A}^{(t+1)}\rangle_{j^{\mathrm{in}}_{n}}$ are corrected as follows:
$\begin{split}&[\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}\\ &\leftarrow\frac{\ell^{({t+1})}_{\mathrm{low},{n}}\sqrt{\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{up},{n}}})}+\ell^{({t+1})}_{\mathrm{up},{n}}\sqrt{\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{low},{n}}})}}{\sqrt{\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{up},{n}}})}+\sqrt{\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{low},{n}}})}}\end{split}$
$\begin{split}&\langle\boldsymbol{A}^{(t+1)}\rangle_{j^{\mathrm{in}}_{n}}\\ &\leftarrow\frac{\ell^{({t+1})}_{\mathrm{up},{n}}-\ell^{({t+1})}_{\mathrm{low},{n}}}{\sigma^{(t+1)}\sqrt{\langle\boldsymbol{C}^{(t+1)}\rangle_{j^{\mathrm{in}}_{n}}}\left(\sqrt{\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{up},{n}}})}+\sqrt{\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{low},{n}}})}\right)}\end{split}$
@@ Appendix C Proof of Proposition 2
(76) $1-\prod_{n=1}^{{N_{\mathrm{in}}}}(1-p^{({t})}_{\mathrm{mut},{n}})\cdot\prod_{n=1}^{{N_{\mathrm{ca}}}}\boldsymbol{q}^{(t)}_{n,1}\enspace,$
@@
Appendix D Additional Experimental Results
Figure 6 shows the average computation time per iteration and its cumulative sum SphereIntCOM with $({N_{\mathrm{co}}},{N_{\mathrm{in}}},{N_{\mathrm{ca}}})=(6,6,6)$. Figure 7 shows the best evaluation values varying the numbers of dimensions as $({N_{\mathrm{co}}},{N_{\mathrm{in}}},{N_{\mathrm{ca}}})=(2,2,2),(4,4,4),(6,6,6)$.

### v6

@@ Appendix A Procedure of Modified Margin Correction
(69) $\exists i\in\{1\!:\!\lambda,\ldots,\mu\!:\!\lambda\}\kern 5.0pt\text{s.t.}\kern 5.0pt[\textsc{Enc}(\boldsymbol{x}_{i})]_{j^{\mathrm{in}}_{n}}\neq[\textsc{Enc}(\boldsymbol{m}^{(t)})]_{j^{\mathrm{in}}_{n}}$
For each $n\in\{1,\ldots,{N_{\mathrm{in}}}\}$, the margin correction updates $\boldsymbol{m}^{(t+1)}{\mskip-2.0mu}]_{j^{\mathrm{in}}_{n}}$ and $\langle\boldsymbol{A}^{(t)}\rangle_{j^{\mathrm{in}}_{n}}$ according to whether $n$ is the dimension in which the successful integer mutation occurred.
Case of $\textnormal{{Enc}}(\boldsymbol{m}^{(t+1)})]_{j^{\mathrm{in}}_{n}}\in\{z_{n,1},z_{n,L_{n}}\}$.
@@
(71) $p^{({t+1})}_{\mathrm{mut},{n}}\leftarrow\max\left\{\alpha,p^{({t+1})}_{\mathrm{mut},{n}}\right\}\kern 5.0pt$.
@@
(72) $p^{({t+1})}_{\mathrm{mut},{n}}\leftarrow\max\left\{\alpha,\min\left\{p^{({t+1})}_{\mathrm{mut},{n}},p^{({t})}_{\mathrm{mut},{n}}\right\}\right\}\kern 5.0pt$.
@@
Finally, we correct $\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}$ as follows:
@@
Case of $\textnormal{{Enc}}(\boldsymbol{m}^{(t+1)})]_{j^{\mathrm{in}}_{n}}\in\{z_{n,2},\ldots,z_{n,L_{n}-1}\}$.
Let us denote the nearest two thresholds to $\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}$ as follows:
@@
(75) $p^{({t+1})}_{\mathrm{low},{n}}$ $=\Pr\left([\boldsymbol{v}]_{j^{\mathrm{in}}_{n}}\leq\ell^{({t+1})}_{\mathrm{low},{n}}\right)\kern 5.0pt$,
(76) $p^{({t+1})}_{\mathrm{up},{n}}$ $=\Pr\left(\ell^{({t+1})}_{\mathrm{up},{n}}<[\boldsymbol{v}]_{j^{\mathrm{in}}_{n}}\right)\kern 5.0pt,$
(77) $p^{({t+1})}_{\mathrm{mid},{n}}$ $=1-p^{({t+1})}_{\mathrm{low},{n}}-p^{({t+1})}_{\mathrm{up},{n}}\kern 5.0pt$.
@@
(89) $p^{({t+1})}_{\mathrm{mut},{n}}\leftarrow p^{({t+1})}_{\mathrm{low},{n}}+p^{({t+1})}_{\mathrm{up},{n}}\kern 5.0pt$.
Finally, $\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}$ and $\langle\boldsymbol{A}^{(t+1)}\rangle_{j^{\mathrm{in}}_{n}}$ are corrected as follows:
$[\boldsymbol{m}^{(t+1)}]_{j^{\mathrm{in}}_{n}}$ $\leftarrow\frac{\ell^{({t+1})}_{\mathrm{low},{n}}\sqrt{\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{up},{n}}})}+\ell^{({t+1})}_{\mathrm{up},{n}}\sqrt{\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{low},{n}}})}}{\sqrt{\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{up},{n}}})}+\sqrt{\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{low},{n}}})}}$
$\langle\boldsymbol{A}^{(t+1)}\rangle_{j^{\mathrm{in}}_{n}}$ $\leftarrow\frac{\ell^{({t+1})}_{\mathrm{up},{n}}-\ell^{({t+1})}_{\mathrm{low},{n}}}{\sigma^{(t+1)}\sqrt{\langle\boldsymbol{C}^{(t+1)}\rangle_{j^{\mathrm{in}}_{n}}}\left(\sqrt{\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{up},{n}}})}+\sqrt{\chi^{2}_{\mathrm{ppf}}({1-2p^{({t+1})}_{\mathrm{low},{n}}})}\right)}$
@@ Appendix C Proof of Proposition 2
(93) $1-\prod_{n=1}^{{N_{\mathrm{in}}}}(1-p^{({t})}_{\mathrm{mut},{n}})\cdot\prod_{n=1}^{{N_{\mathrm{ca}}}}\boldsymbol{q}^{(t)}_{n,1}\kern 5.0pt$,